OpenStudy (anonymous):

What is the slope of the line passing through the point (4, -2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction

OpenStudy (radar):

To answer this question you only need to look at the line y=4x+2

OpenStudy (radar):

What is the slope of this line y=4x+2?

OpenStudy (radar):

When you have the slope, take the negative reciprocal and you will have the slope if the line thaT is perpendicular and the answer.

OpenStudy (radar):

Reciprocal you say, what is that..... A reciprocal of a number is another number or fraction that when multiplied of the number equals 1 example what is the reciprocal of 2? It would be 1/2 as 2 * 1/2 = 1.

OpenStudy (radar):

Have you figured out what is the slope of this line?: y=4x + 2

OpenStudy (anonymous):

4?

OpenStudy (radar):

Yes, and what is its reciprocal?

OpenStudy (anonymous):

1

OpenStudy (radar):

If i multiply 1 times 4 will the answer be 1? remember the reciprocal when multiplied by the number (in this case 4) will be 1. Like 4x=1 what is x?

OpenStudy (anonymous):

4

OpenStudy (radar):

if x is 4 then 4x= 4*4=16 which is not 1. What times 4 equals 1?

OpenStudy (anonymous):

nothing

OpenStudy (radar):

4 times nothing is nothing that is not a 1 4x=1 solve for x

OpenStudy (radar):

4x=1 solve for x divide both sides by 4 x=1/4 Note that 4 * 1/4 = 1 Now make it negative by multiplying by -1 or -1/4 is the slope of the required line.